# Understanding Chain-of-Thought in LLMs through Information Theory

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive performance in complex reasoning tasks through Chain-of-Thought (CoT) reasoning, allowing models to break down problems into manageable sub-tasks. However, existing CoT evaluation techniques either require annotated CoT data or fall short in accurately assessing intermediate reasoning steps, leading to high rates of false positives. In this paper, we formalize CoT reasoning in LLMs through an information-theoretic lens. Specifically, our framework quantifies the ‘information gain’ at each reasoning step, enabling the identification of failure modes in LLMs without the need for expensive annotated datasets. We demonstrate the efficacy of our approach through extensive experiments on toy and GSM-8K data, where it significantly outperforms existing outcome-based methods by providing more accurate insights into model performance on individual tasks.
Denotes equal contribution, where ordering was determined through a coin flip.
Corresponding authors: jeanfrancois@bytedance.com and faaiz.taufiq@bytedance.com.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, from complex reasoning to code generation (Chowdhery et al., 2024; OpenAI et al., 2024; Bubeck et al., 2023; Anil et al., 2023). Many of these advances can be attributed to Chain-of-Thought (CoT) reasoning (Wei et al., 2024; Nye et al., 2021; Li et al., 2024), which involves breaking down complex problems into a series of intermediate steps, mirroring human-like reasoning processes. The success of CoT reasoning, particularly in domains such as mathematics, logic, and multi-step decision-making, has led researchers and developers to incorporate CoT-like features directly into model training, i.e. the FLAN family of models (Chung et al., 2022; Wei et al., 2022).
@@
Previous work in this area has proposed “Process Supervision(Lightman et al., 2023), which requires expensive, human-annotated step-by-step data. While effective, this approach is often impractical due to the high cost and effort of creating large-scale annotated datasets. In turn, alternative methods have recently been proposed, such as outcome reward modelling (Havrilla et al., 2024) or the Math-Shepherd (Wang et al., 2024b). Both these approaches avoid reliance on annotated step-wise CoT data by instead modelling the correctness of each step based on the correctness of final outputs. However, as we demonstrate in this paper, these methods can be unsound for detecting incorrect reasoning steps and can thus lead to a high false-positive rate in certain scenarios.
To address these shortcomings, we employ an information-theoretic approach, grounded in the following key insight: Each correct step in a reasoning process should provide valuable and relevant information that aids in predicting the final correct outcome. Building on this insight, we develop a framework to quantify the “information gain” after each sub-task in the reasoning process, without the need for step-by-step annotations. This enables us to detect sub-tasks that fail to contribute meaningful information toward the correct solution, signalling potential errors or irrelevant steps in the model’s reasoning. In addition, we also introduce a practical algorithm to assess LLM performance across various sub-tasks within a Chain-of-Thought (CoT) reasoning process.
The key contributions of this paper are as follows:
@@
    We develop a framework for sequential applications of sub-tasks, e.g. Chain-of-Thought and provide a rigorous language to describe and detect detect failure modes in LLMs.
@@
    We validate our methods on extensive toy data and the GSM-8K dataset (Cobbe et al., 2021). Our method effectively identifies failure modes in CoT reasoning, unlike baselines like outcome reward modelling (Havrilla et al., 2024) and Math-Shepherd (Wang et al., 2024b), which rely on final accuracy and tend to increase false positives in error detection.
@@ 2 Proposed Framework: Setup and Notation
Before diving into our framework, we first provide a high-level overview and notation on how LLM generation will be treated throughout this paper. This will allow us to set the foundation for describing our information-theoretic framework. In particular, following the approach in González and Nori (2023), we view LLMs as abstract execution machines with a natural language interface. From this perspective, prompts are designed to solve specific problems (e.g., mathematical or logical problems), and the LLM processes the information in the prompt to generate an output.
We now define the notation for a typical prompt as a combination of two components:
@@
    An initial state, represented by a random variable $X_{0}\in\mathcal{X}$, denotes information provided in the prompt that the LLM must operate on to obtain the queried information.
@@
    A task $\lambda\in\Upsilon$ (e.g., addition followed by multiplication) which encapsulates how the LLM should process information in $X_{0}$.
Given the prompt, defined as a tuple $(X_{0},\lambda)$, the state $X_{1}$ represents the result of applying task $\lambda$ to the initial state $X_{0}$. Formally, we denote this using the update mapping $\Lambda:\mathcal{X}\times\Upsilon\rightarrow\mathcal{X}$ which outputs the updated state $X_{1}$ by applying the task $\lambda$ on $X_{0}$, i.e. $X_{1}=\Lambda(X_{0},\lambda)$. This updated state is then used to obtain the final output, denoted by $Y\in\mathcal{X}$, by extracting only the information in $X_{1}$ which is relevant to the queried final answer. This notation defines a prompt that instructs a model to process information drawn from some initial distribution $p(X_{0})$ (e.g., mathematical problems).
Let us use the following simple example to illustrate the notation:
Prompt:“James has 3 apples and Abbey has 9. How many apples do the two have in total?”
Here, using the above notation, the initial state $x_{0}$ denotes the information “James has 3 apples; Abbey has 9 apples”, and $\lambda$ denotes the addition task. Next, $x_{1}=\Lambda(x_{0},\lambda)$ represents the updated information after correctly performing the addition operation, i.e. $x_{1}=$“James has 3 apples; Abbey has 9 apples; The two have 12 apples in total”. The final output, $y$, is then obtained by simply extracting the total number of apples from $x_{1}$, i.e. “The two have 12 apples in total” [^1] Our setup also encapsulates cases with ambiguous (or multiple correct) responses for a given task $\lambda$ and initial state $x_{0}$. In this case, $\Lambda(x_{0},\lambda)$ is a random variable with distribution $p(X_{1}\mid X_{0}=x_{0})$. Therefore, for generality, we treat $\Lambda(x_{0},\lambda)$ as a random variable from now on.. With this basic notation established, we now consider compositions of tasks, enabling us to formalize the Chain of Thought (CoT) process in LLMs.
@@ 2.1 Compositionality
Many mathematical or logical problems require a sequential application of operations. Our notation is also amenable to such problems as it accommodates the composition of tasks. Consider a problem which requires two successive steps to arrive at the correct output:
Prompt: “Solve for z=2×(x+y) where x = 12 and y = $\textbf{Prompt: }\textit{``Solve for $z=2\times(x+y)$ where x = 12 and y = 13''}.$” . (1)
In this example, first, we apply the addition operation to find the value of $x+y$, and next, we apply the multiplication operation to find the value of $z$. Using our notation this can be expressed as $\Lambda(x_{0},\lambda_{1}\circ\lambda_{2})$, where $\lambda_{1},\lambda_{2}$ denote the addition and multiplication tasks respectively. The following property allows us to concretely define the application of compositional task $\lambda_{1}\circ\lambda_{2}$:
@@
We say that an update rule $\Lambda:\mathcal{X}\times\Upsilon\rightarrow\mathcal{X}$ is compositionally consistent if:
@@
Going back to the prompt in (1), Figure 1 shows that the model first computes $x+y$, and next multiplies the result by $2$.
Figure 1: Solving the problem in prompt (1) requires compositional application of tasks.
Here, we refer to $X_{1},X_{2}$ as intermediate states and $Y$ is the correct final output. More generally, if a problem statement requires sequential application of $T$ sub-tasks, $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{T}$, then the Chain-of-Thought (CoT) reasoning is divided up into $T$ steps, where the output of the $t$’th step is recursively defined as $X_{t}=\Lambda(X_{t-1},\lambda_{t})$ for $t\in\{1,\dots,T\}$. Finally, the overall true output $Y$ is obtained by extracting the queried information from the final state $X_{T}$.
Having established a formal language for the sequential application of tasks, e.g. CoT, we now turn towards how a task may be divided into such a sequence of intermediate sub-tasks.
@@ 2.2 Primitive tasks
Note that the decomposition is not unique but the set of components is. In some cases, there may exist distinct permutations of primitive tasks which compose to yield the same task as is common in many associative operations. As an example, in the context of mathematical problem-solving, the basic arithmetic operation could be considered primitive. The composition of these primitive tasks allows us to construct extremely complex operations. Just like in linear algebra, we define the span of these tasks as the set obtained by their sequential applications.
@@
Let $\Phi\subseteq\Upsilon$ be a set of tasks, then:
@@
The set $\textup{Span}(\Phi)$ comprises all the tasks that can be applied by composing sub-tasks in the set $\Phi$. This means that any compositionally consistent update rule $\Lambda$ which is well-defined on the set of tasks $\Phi$ will also be well-defined on $\textup{Span}(\Phi)$. However, this $\Lambda$ may still be ill-defined for any task not in this span. This limitation is captured by the concept of unidentifiability, which plays a central role in determining the boundaries of what a model can and cannot infer.
@@ 2.3 Unidentifiability
The unidentifiability of tasks forms a key part of our framework. It directly addresses the fundamental challenge that models, such as LLMs, face when dealing with unseen tasks. If a task $\lambda$ lies outside of $\textup{Span}(\Phi)$, the span of tasks the model has been trained on, then the model cannot be expected to infer or apply it correctly. In other words, the model’s capacity is constrained by the identifiability of tasks within the training set. This notion and formalization of unidentifiability allows us to highlight a critical limitation in the generalization of models: tasks not encountered during training cannot be reliably executed, as they remain beyond the model’s learned task span. More formally:
@@
Let $\Phi\subseteq\Upsilon$ be any set of tasks, then a tasks $\lambda$ is said to be unidentifiable in $\Phi$ iff, $\lambda\not\in\textup{Span}(\Phi)$.
[…]
With this general framework in place, we can now turn this theoretical foundation into a practical algorithm for detecting unidentifiable sub-tasks. Specifically, we explore how the notion of unidentifiability can be combined with information-theoretic approaches to detect failure points in LLMs.
@@ 3.1 An information-theoretic perspective
To make the concept of unidentifiability practical in the context of CoT generations, we begin by introducing the fundamental assumption. The core assumption in our approach is that each correctly executed CoT reasoning step should contribute meaningful and relevant information that aids in predicting the correct final output, denoted as $Y$. If we encounter a step after which the amount of information regarding $Y$ stops increasing, then we can take this as an indication of an incorrectly executed task. We concretise this assumption using using our notation from the previous section:
@@
$\lambda=\lambda_{1}\circ\dots\lambda_{k-1}\circ\lambda_{k}\circ\dots\circ\lambda_{T}\quad\text{ and }\quad\lambda^{\prime}=\lambda_{1}\circ\dots\lambda_{k-1}\circ\lambda^{\prime}_{k}\circ\dots\circ\lambda^{\prime}_{T^{\prime}},$
where $\lambda^{\prime}_{k}$ is unidentifiable in $\{\lambda_{1},\dots,\lambda_{T}\}$. Then, the intermediate states corresponding to the tasks $\lambda,\lambda^{\prime}$ have the following Bayesian network:
@@
Figure 2: Bayesian network
@@
The Bayesian network in Figure 2 implies that for any two reasoning paths which diverge at step $k$, the future states $X_{i}$ and $X^{\prime}_{j}$ for any $i,j\geq k$ satisfy the conditional independence $X_{i}\perp\!\!\!\perp X^{\prime}_{j}\mid X_{k-1}$. Consequently, once we apply $\lambda^{\prime}_{k}$, the subsequent states along the new reasoning path (in red) add no information regarding the subsequent states or the output of the original path (in green). Hence the figure represents the fact that, for any given input, the output of $\lambda_{k}$ (top fork) contains no information regarding the output of any other primitive task $\lambda^{\prime}_{k}$ (bottom fork).
Now that we have formalised our key information-theoretic assumption on the ground-truth CoT process, we turn towards the model behaviour on unidentifiable tasks in the following section.
@@ 3.2 Task execution in LLMs
To operationalise our framework, we formally distinguish between the model i.e. LLM’s task execution and the ground truth process which arises from following the instructions correctly. To this end, we explicitly define how an LLM interprets a specified task $\lambda$ using the update rule, $\Lambda^{M}(X_{0},\lambda)$, which is in general distinct from the ground truth update rule $\Lambda(X_{0},\lambda)$.
Here, one option would be to consider the idealised setting where the model learns to perfectly follow some of the primitive tasks available in the training data. However, this may be considered too restrictive since in reality most LLMs do not always follow a “learned” task perfectly. Instead, we consider a much weaker assumption that the model cannot correctly execute a task which is unidentifiable in the training data. To this end, suppose $\Gamma^{M}\subseteq\Gamma$ denotes the primitive tasks available in the LLM training data. Concretely, we make the following assumption on LLM’s task execution.
@@
Assumption 3.2 means that for any task which we would like the LLM to apply, the LLM ends up executing some task in $\textup{Span}(\Gamma^{M})$ which the model has been trained on. In other words, the model’s execution is restricted only to the tasks which could be inferred from the training data (i.e. in $\textup{Span}(\Gamma^{M})$). Moreover, this assumption also allows us to encapsulate cases where the model does not follow the correct instructions or does not decompose a given task correctly.
@@
Moreover, we use $Y^{M}$ to denote the model’s final output. Next, using this notation, we present the conditional independence which must hold if the model encounters an unidentifiable intermediate task along its CoT reasoning path.
@@
Theorem 3.3 shows that under Assumptions 3.1 and 3.2, when the model encounters an unidentifiable task (i.e. $\lambda_{k}$ in Theorem 3.3) in its Chain-of-Thought reasoning, the model output satisfies the conditional independence in Equation (2). More concretely, after a model’s CoT reasoning diverges from the ground truth at step $k$, every subsequent step adds no additional information regarding the correct final output $Y$. In practice, this ‘information’ is measured by checking if the model’s confidence about the final output $Y$ increases after each step. This is formalised in the next section.
3.3 Testing for unidentifiability using information gain
Having established all the essential components of our framework, we can now provide a concrete description of how to practically identify unidentifiable sub-tasks using information theory. As is common in the literature (Wang et al., 2024b; Havrilla et al., 2024), we assume access to a dataset consisting of prompts and their corresponding final answers, obtained by correctly applying the task $\lambda$. This dataset is denoted as $\mathcal{D}_{\lambda}\coloneqq\{(x_{0}^{i},y^{i})\}_{i=1}^{n}$.
Additionally, recall that $X^{M}_{j}$ and $X^{M}_{j-1}$ represent the model’s chain of thought (CoT) reasoning at steps $j$ and $j-1$, respectively. Consequently, each element in the conditional independence statement in Equation (2) can be derived from the data and/or the model.
To this end, we consider the mutual information between $Y$ and $X^{M}_{j}$ conditional on $X^{M}_{j-1}$, denoted by $\mathcal{I}(Y;X^{M}_{j}\mid X^{M}_{j-1})$. This conditional mutual information term intuitively represents the additional information contributed by the $j$’th step of CoT, that is relevant for predicting the ground truth final output $Y$. Therefore, we refer to $\mathcal{I}(Y;X^{M}_{j}\mid X^{M}_{j-1})$ as the information gain at step $j$.
It follows from Theorem 3.3 that if an LLM encounters a sub-task at step $i$ which is unidentifiable in its training data, no subsequent step should contribute any additional information relevant for predicting $Y$ (i.e. the information gain should remain 0 after step $i$). If, on the other hand, we observe that $\mathcal{I}\left(Y;X^{M}_{j}\mid X^{M}_{j-1}\right)>0$ for some $j\geq i$, then under Assumptions 3.1 and 3.2, the task $\lambda_{i}$ is not unidentifiable. To estimate the information gain in practice, we use the following result:

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive performance in complex reasoning tasks through the use of Chain-of-Thought (CoT) reasoning, allowing models to break down problems into manageable sub-tasks. However, existing CoT evaluation techniques either require annotated CoT data or fall short in accurately assessing intermediate reasoning steps, leading to high rates of false positives. In this paper, we formalize CoT reasoning in LLMs through an information-theoretic lens. Specifically, our framework quantifies the ‘information-gain’ at each reasoning step, enabling the identification of failure modes in LLMs without the need for expensive annotated datasets. We demonstrate the efficacy of our approach through extensive experiments on toy arithmetic, GSM8K and PRM800k datasets, where it significantly outperforms existing outcome-based methods by providing more accurate insights into model performance on individual subtasks.
[…]
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, from complex reasoning to code generation [4, 23, 3, 2]. Many of these advances can be attributed to Chain-of-Thought (CoT) reasoning [32, 22, 17], which involves breaking down complex problems into a series of intermediate steps, mirroring human-like reasoning processes. The success of CoT reasoning, particularly in domains such as mathematics, logic, and multi-step decision-making, has led researchers to incorporate CoT-like features directly into model training, i.e. the FLAN family of models [5, 31].
@@
Previous work in this area has proposed “Process Supervision[18], which requires expensive, human-annotated step-by-step data. While effective, this approach is often impractical due to the high cost and effort of creating large-scale annotated datasets. In turn, alternative methods have recently been proposed, such as outcome reward modelling [11] or the Math-Shepherd [29]. Both these approaches avoid reliance on costly annotated step-wise CoT data by instead modelling the correctness of each step based on the correctness of final outputs. However, as we show later, these methods can be unsound for detecting incorrect reasoning steps and can thus lead to a high false-positive rate in certain scenarios.
To address these shortcomings, we employ an information-theoretic approach, grounded in the following key insight: Each correct step in a reasoning process should provide valuable and relevant information that aids in predicting the final correct outcome. Building on this insight, we develop a framework to quantify the “information-gain” after each sub-task in the reasoning process, without the need for step-by-step annotations. This enables us to detect sub-tasks that fail to contribute meaningful information toward the correct solution, signalling potential errors or irrelevant steps in the model’s reasoning. In addition, we also introduce a practical algorithm to assess LLM performance across various sub-tasks within a Chain-of-Thought (CoT) reasoning process. The key contributions of this paper are as follows:
@@
    We develop a framework for sequential applications of sub-tasks, e.g. Chain-of-Thought and provide a rigorous language to identify failure modes in LLMs.
@@
    We validate our methods on extensive toy data, the GSM8K [6] as well as the PRM800K [18] dataset. Our method effectively identifies failure modes in CoT reasoning, unlike baselines like outcome reward modelling [11] and Math-Shepherd [29], which rely on final accuracy and tend to increase false positives in error detection.
@@ 2 Proposed Framework: Setup and Notation
Before diving into our framework, we first provide a high-level overview and notation on how LLM generation will be treated throughout this paper. This will allow us to set the foundation for describing our information-theoretic framework. In particular, following the approach in [10], we view LLMs as abstract execution machines with a natural language interface. From this perspective, prompts are designed to solve specific problems (e.g., mathematical or logical problems), and the LLM processes the information in the prompt to generate an output.
We define a typical prompt as a combination of two parts:
@@
    An initial state, represented by the random variable $X_{0}\in\mathcal{X}$, encapsulates the prompt-provided information that the LLM processes to derive the queried result.
@@
    A task $\lambda\in\Upsilon$ (e.g., addition then multiplication) defines how the LLM processes $X_{0}$.
Given the prompt, defined as a tuple $(X_{0},\lambda)$, the state $X_{1}$ represents the result of applying task $\lambda$ to the initial state $X_{0}$. Formally, we denote this using the update mapping $\Lambda:\mathcal{X}\times\Upsilon\rightarrow\mathcal{X}$ which outputs the updated state $X_{1}$ by applying the task $\lambda$ on $X_{0}$, i.e. $X_{1}=\Lambda(X_{0},\lambda)$. This updated state is then used to obtain the final output, denoted by $Y\in\mathcal{X}$, by extracting only the information in $X_{1}$ which is relevant to the queried final answer. This notation defines a prompt that instructs a model to process information drawn from some initial distribution $p(X_{0})$ (e.g., math problems).
We use a simple example to illustrate this notation:
Prompt: “Solve for $z=2\,(u+v)$, $u=12$, $v=13$ (1)
Here, the initial state $x_{0}$ denotes the information $u$ = 12, $v$ = 13”, and $\lambda$ denotes the task of finding $z$ (i.e. addition followed by multiplication). Next, $x_{1}=\Lambda(x_{0},\lambda)$ represents the updated information after correctly performing the addition operation, i.e. $x_{1}$ is $u$ = 12, $v$ = 13 and $z$ = 50”. The final output, $y$, is then obtained by simply extracting the value of $z$ from $x_{1}$, i.e. $z$ = 50”.
Remark Our setup also encapsulates cases with ambiguous (or multiple correct) responses for a given task $\lambda$. In this case, $\Lambda(x_{0},\lambda)$ is a random variable with distribution $p(X_{1}\mid X_{0}=x_{0})$. Therefore, for generality, we treat $\Lambda(x_{0},\lambda)$ as a random variable from now on.
@@ 2.1 Compositionality
Many mathematical or logical problems, such as the one in (1), require sequential application of operations. Our notation is also amenable to such problems as it accommodates the composition of tasks.
For example, one way to address prompt (1) involves first adding $u$ and $v$, and next, multiplying the result by $2$ to find $z$. Using our notation, this can be expressed as $\Lambda(x_{0},\lambda_{1}\circ\lambda_{2})$, where $\lambda_{1},\lambda_{2}$ denote the addition and multiplication tasks respectively. The following property allows us to define the application of compositional task $\lambda_{1}\circ\lambda_{2}$:
@@
Returning to the prompt in (1), Figure 1 shows that the model first computes $u+v$, then multiplies the result by $2$.
Figure 1: Prompt (1) requires compositional application of tasks.
Here, we refer to $X_{1},X_{2}$ as intermediate states and $Y$ is the correct final output. In general, if a problem statement requires sequential application of $T$ sub-tasks, $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{T}$, then the Chain-of-Thought (CoT) reasoning is divided up into $T$ steps, where the $t$’th step is recursively defined as $X_{t}=\Lambda(X_{t-1},\lambda_{t})$ for $t\in\{1,\dots,T\}$. Finally, the overall true output $Y$ is obtained by extracting the queried information from the final state $X_{T}$.
Having established a formal language for the sequential application of tasks, we now turn towards how a task may be divided into such a sequence of intermediate sub-tasks.
@@ 2.2 Primitive Tasks
Note that the decomposition is not unique but the set of components is. In some cases, there may exist distinct permutations of primitive tasks which compose to yield the same task as is common in many associative operations. As an example, in the context of mathematical problem-solving, the basic arithmetic operations could be considered primitive. The composition of these primitive tasks allows us to construct extremely complex operations. Just like in linear algebra, we define the span of these tasks as the set obtained by their sequential applications.
@@
Let $\Phi\subseteq\Upsilon$ be a set of tasks:
@@
The set $\textup{Span}(\Phi)$ comprises all the tasks that can be applied by composing sub-tasks in the set $\Phi$. This means that any compositionally consistent update rule $\Lambda$ which is well-defined on the set of tasks $\Phi$ will also be well-defined on $\textup{Span}(\Phi)$. However, this $\Lambda$ may still be ill-defined for any task not in this span. This limitation, known as unidentifiability, defines the boundaries of a model’s inferences.
@@ 2.3 Unidentifiability
The unidentifiability of tasks forms a key part of our framework. It directly addresses the fundamental challenge that models, such as LLMs, face when dealing with unseen tasks. If a task $\lambda$ lies outside of $\textup{Span}(\Phi)$, the span of tasks the model has been trained on, then the model cannot be expected to infer or apply it correctly. In other words, the model’s capacity is constrained by the identifiability of tasks within the training set. This notion and formalization of unidentifiability allows us to highlight a critical limitation in the generalization of models: tasks not encountered during training cannot be reliably executed, as they remain beyond the model’s learned task-span. More formally:
@@
A task $\lambda$ is unidentifiable in a set $\Phi\subseteq\Upsilon$ if and only if $\lambda\not\in\textup{Span}(\Phi)$.
Remark   In practice, unidentifiability may depend on the initial state $X_{0}$, i.e. an LLM might accurately perform addition for 2-digit numbers but fail with 10-digit numbers [24]. For more details, see Appendix 8.1.
Building on this framework, we propose an algorithm that integrates unidentifiability with information-theoretic methods to detect CoT reasoning failures.
@@ 3.1 An information-theoretic perspective
To apply unidentifiability in CoT generations, we introduce a fundamental assumption: each correct CoT step should provide meaningful information aiding the prediction of $Y$. If a step ceases to increase information about $Y$, it indicates an incorrect execution. We formalize this assumption using our notation from the previous section:
@@
$\lambda$ $=\lambda_{1}\circ\dots\lambda_{k-1}\circ\lambda_{k}\circ\dots\circ\lambda_{T}\quad\text{ and }$
$\lambda^{\prime}$ $=\lambda_{1}\circ\dots\lambda_{k-1}\circ\lambda^{\prime}_{k}\circ\dots\circ\lambda^{\prime}_{T^{\prime}},$
where $\lambda^{\prime}_{k}$ is unidentifiable in $\{\lambda_{1},\dots,\lambda_{T}\}$. Then, the intermediate states corresponding to the tasks $\lambda,\lambda^{\prime}$ have the Bayesian network in Figure 2.
@@
Figure 2: Bayesian network corresponding to Assumption 3.1.
@@
The Bayesian network in Figure 2 implies that if we encounter an unidentifiable task ($\lambda^{\prime}_{k}$) at step $k$ of the reasoning path, the future states $X_{i}$ and $X^{\prime}_{j}$ for any $i,j\geq k$ satisfy the conditional independence $X_{i}\perp\!\!\!\perp X^{\prime}_{j}\mid X_{k-1}$. Consequently, once we apply $\lambda^{\prime}_{k}$, the subsequent states along the new reasoning path (in red) add no information regarding the subsequent states or the output of the original path (in green). Hence the figure represents the fact that, for any given input, the output of $\lambda_{k}$ (top fork) contains no information regarding the output of any other primitive task $\lambda^{\prime}_{k}$ (bottom fork).
With our key assumption on the ground-truth CoT process formalized, we now consider the model’s CoT behaviour.
@@ 3.2 Task execution in LLMs
To operationalise our framework, we formally distinguish between the model i.e. LLM’s task execution and the ground truth process which arises from following the instructions correctly. To this end, we explicitly define how an LLM interprets a specified task $\lambda$ using $\Lambda^{M}(X_{0},\lambda)$, which is in general distinct from the ground truth update rule $\Lambda(X_{0},\lambda)$.
Here, one option would be to consider the idealised setting where the model learns to perfectly follow some of the primitive tasks available in the training data. However, this may be considered too restrictive since in reality most LLMs do not always follow a “learned” task perfectly. Instead, we consider a much weaker assumption that the model cannot correctly execute a task which is unidentifiable in the training data. Concretely, suppose $\Gamma^{M}\subseteq\Gamma$ denotes the primitive tasks available in the LLM training data. Then, we make the following assumption on LLM’s task execution.
@@
Assumption 3.2 means that for any task which we would like the LLM to apply, the LLM ends up executing some task in $\textup{Span}(\Gamma^{M})$ which the model has been trained on. In other words, the model’s execution is restricted only to the tasks which could be inferred from the training data (i.e. in $\textup{Span}(\Gamma^{M})$). Moreover, this assumption also allows us to encapsulate cases where the model does not follow correct instructions or does not decompose a task correctly.
@@
Moreover, we use $Y^{M}$ to denote the model’s final output. Next, using this notation, we present the conditional independence which must hold if the model encounters an unidentifiable intermediate task along its reasoning path.
@@
Theorem 3.3 shows that under Assumptions 3.1 and 3.2, when the model encounters an unidentifiable task (i.e. $\lambda_{k}$ in Theorem 3.3) in its Chain-of-Thought reasoning, the model output satisfies the conditional independence in Equation (2). In practice, this means that if at step $k$, a model encounters a reasoning step which is necessary for obtaining the correct answer and is unidentifiable in the training data, the CoT reasoning diverges from the ground truth at this step and every subsequent step adds no additional information regarding the correct final output Y. This ‘information’ can be measured by checking if the model’s confidence about the final output $Y$ increases after each step. This is formalised in the next section.
3.3 Testing for unidentifiability using information-gain
With our framework established, we now describe how to detect unidentifiable sub-tasks using information theory. Following common practice [29, 11], we assume access to a dataset of prompts and correct final answers, derived by applying task $\lambda$, denoted as $\mathcal{D}_{\lambda}\coloneqq\{(x_{0}^{i},y^{i})\}_{i=1}^{n}$. Recall that $X^{M}_{j}$ and $X^{M}_{j-1}$ represent the model’s chain of thought (CoT) reasoning at steps $j$ and $j-1$, respectively. Consequently, each element in the conditional independence statement in Equation (2) can be derived from the data and/or the model.
To this end, we consider the mutual information between $Y$ and $X^{M}_{j}$ conditional on $X^{M}_{j-1}$, denoted by $\mathcal{I}(Y;X^{M}_{j}\mid X^{M}_{j-1})$. This conditional mutual information term intuitively represents the additional information contributed by the $j$’th step of CoT, which is relevant for predicting the ground truth final output $Y$. Therefore, we refer to $\mathcal{I}(Y;X^{M}_{j}\mid X^{M}_{j-1})$ as the information-gain at step $j$.
It follows from Theorem 3.3 that if an LLM encounters a sub-task at step $i$ which is unidentifiable in its training data, no subsequent step should contribute any additional information relevant for predicting $Y$ (i.e. the information-gain should remain 0 after step $i$). If, on the other hand, we observe that $\mathcal{I}\left(Y;X^{M}_{j}\mid X^{M}_{j-1}\right)>0$ for some $j\geq i$, then under Assumptions 3.1 and 3.2, the task $\lambda_{i}$ is not unidentifiable. To estimate the information-gain in practice, we use the following result: